Chapter 13, Section 13.1, Problem 021f The following table gives information on the amount of sugar (in grams) and the calorie count in one serving of a sample of 13 varieties of Kellogg's cereal. Sugar (grams) 6 15 12 11 8 6 7 2 7 14 20 3 13 Calories 120 200 140 110 130 80 210 100 120 190 190 110 120 The predictive regression equation of the number of calories on the amount of sugar is y ^ = 91.296 + 5.106 x , where x is amount of sugar (in grams) and y is calories. Estimate the calorie count for a cereal with 52 grams of sugar per serving. Round your answer to the nearest integer. calories Comment on this finding. the tolerance is +/-2%
Answer #1

Given that, the regression equation is,

у = 91.296 5. 1062.

We want to predict y for the value of x = 52 grams is,

\hat y = 91.296 + (5.106*52) = 91.296 + 265.512 = 356.808

\hat y \approx 357

Therefore, the calorie count for a cereal with 52 grams of suger per serving is 357